# Symmetric Linear Bandits with Hidden Symmetry

### v1

@@ 6 Conclusion and Future Work
In this paper, we study symmetric linear stochastic bandits in high dimensions, where the linear reward function is invariant with respect to some hidden subgroup $\mathcal{G}\leq\mathcal{S}_{d}$. We first prove that no algorithm can gain any advantage solely by knowing $\mathcal{G}\leq S_{d}$. Given this, we introduce a cardinality condition on the hidden subgroup $\mathcal{G}$, allowing the learner to overcome the curse of dimensionality. Under this condition, we propose novel model selection algorithms that achieve regrets of $\tilde{O}(d_{0}^{1/3}T^{2/3})$ and $\tilde{O}(d_{0}\sqrt{T})$ with an additional assumption on the well-separated partition. For future work, we will explore convex relaxation techniques for efficient computation, leveraging specific structures of symmetries. In addition, we will investigate whether there exists an adaptive algorithm that achieves $O(d_{0}\sqrt{T})$ in the case of a well-separatedness, while also achieving $O(d_{0}^{1/3}T^{2/3})$ in the worst case.
@@ Appendix A Impossibility Result of Learning with General Hidden Subgroups
Now, $\Theta$ can be used as a family of problem instances for minimax lower bound of linear bandit. By Theorem 24.2 in [24], if $T\geq d/2$, we have that with the choice of $\varepsilon=\frac{1}{4\sqrt{3}}\sqrt{d/T}$, one has that
$\mathbf{R}_{T}\geq\frac{16}{\sqrt{3}}d\sqrt{T}.$ (18)
@@ Appendix B The Case of Hidden Subgroups with Subexponential Size
B.1 High Probability Prediction Error of Model Selection
@@ B.2 Regret Upper bound
First, let consider Proposition 15, with fixed constant $\delta_{\,\overline{\!{\mathcal{M}}}}(X)=1/2$, and recall that $\log(M)$ and $\log(\,\overline{\!{M}})$ are both $O(d_{0}\log(d))$. Therefore, for the choice of $t_{1}=\Omega(d_{0}\log(d/\delta))$, with probability at least $21-\delta/$ 2, one has that
$\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}^{2}\leq\frac{2}{t_{1}}\left\|X(\theta_{\star}-\widehat{\theta}_{t_{1}})\right\|_{2}^{2}$ (32)
@@
$\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}\leq c_{2}\sqrt{\frac{\sigma^{2}d_{0}\log\left(\frac{d}{\delta}\right)}{t_{1}}},$ (34)
@@
Let define the pesudo regret as $\widehat{\mathbf{R}}_{T}=\sum_{t=1}^{T}\left<x_{\star}-x_{t},\theta_{\star}\right>$. We start by simple regret decomposition as follows.
@@
$\leq 2R_{\mathrm{\max}}t_{1}+\sum_{t=t_{1}+1}^{T}\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}\left\|x_{t}-x_{\star}\right\|_{2}$
$\leq 2R_{\mathrm{\max}}t_{1}+\sum_{t=t_{1}+1}^{T}2\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}.$
Now, we invoke Lemma 11. Let $\mathcal{E}$ is the event in that the exploration error is bounded as in Lemma Lemma 11, one has that
$\widehat{\mathbf{R}}_{T}$ $\leq 2R_{\mathrm{\max}}t_{1}+\mathbb{E}\left[\sum_{t=t_{1}+1}^{T}2\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}\bigg{|}\mathcal{E}\right]+T\mathrm{Pr}(\mathcal{E})R_{\mathrm{\max}}$ (36)
$\leq 2R_{\mathrm{\max}}t_{1}+c_{1}\sqrt{\frac{\sigma^{2}d_{0}\log(d/\delta)}{t_{1}}}T+T\delta R_{\mathrm{\max}}.$
Let $\delta=1/T$, and $t_{1}=R_{\mathrm{\max}}^{-\frac{2}{3}}T^{\frac{2}{3}}\sigma^{\frac{2}{3}}(d_{0}\log(dT))^{\frac{1}{3}}$, one has that
$\mathbf{R}_{T}=O\left(R_{\mathrm{\max}}^{\frac{1}{3}}T^{\frac{2}{3}}\sigma^{\frac{2}{3}}(d_{0}\log(dT))^{\frac{1}{3}}\right)$ (37)
@@ C.1 Improve Regret Bound with Well-Separated Partitions
Now, by choosing $t_{2}=\Omega\left(\log(M\delta^{-1})\right)$, by Proposition 15, we have that with probability at least $31-\delta/$ 3,
$\left\|\bm{f}_{\star}-\Pi_{S_{m}}(\bm{f}_{\star})\right\|_{2}^{2}=\left\|X(\theta_{\star}-\Pi_{S_{m}}(\theta_{\star}))\right\|_{2}^{2}\geq\frac{t_{2}}{2}\left\|\theta_{\star}-\Pi_{S_{m}}(\theta_{\star})\right\|_{2}^{2}\geq\frac{t_{2}\varepsilon_{0}^{2}}{4}$ (43)
@@
$\left\|\bm{f}_{\star}-\bm{\widehat{f}}_{m}\right\|_{2}^{2}+2\left<\bm{\eta},\bm{f}_{\star}-\bm{\widehat{f}}_{m}\right>\geq\frac{t_{2}\varepsilon_{0}^{2}}{8}-5\left(\sigma^{2}d_{0}+\sigma^{2}(\log(3M\delta^{-1}))\right).$ (44)
@@
$\frac{t_{2}\varepsilon_{0}^{2}}{8}-5\left(\sigma^{2}d_{0}+\sigma^{2}(\log(M\delta^{-1}))\right)\geq 0;$ (45)
@@
$t_{2}=\Omega\left(\frac{\sigma^{2}d_{0}\log(d\delta^{-1})}{\varepsilon_{0}^{2}}\right).$ (46)
See 18
Proof.
The analysis is an immediate consequence of Lemma 17 with the regret bound of OFUL in the exploitation phase and $\delta=1/T$. ∎

### v2

@@ 7 Conclusion and Future Work
In this paper, we study symmetric linear stochastic bandits in high dimensions, where the linear reward function is invariant with respect to some hidden subgroup $\mathcal{G}\leq\mathcal{S}_{d}$. We first prove that no algorithm can gain any advantage solely by knowing $\mathcal{G}\leq S_{d}$. Given this, we introduce a cardinality condition on the hidden subgroup $\mathcal{G}$, allowing the learner to overcome the curse of dimensionality. Under this condition, we propose novel model selection algorithms that achieve regrets of $\tilde{O}(d_{0}^{2/3}T^{2/3})$ and $\tilde{O}(d_{0}\sqrt{T})$ with an additional assumption on the well-separated partition. For future work, we will explore convex relaxation techniques for efficient computation, leveraging specific structures of symmetries.
@@ Appendix A Impossibility Result of Learning with General Hidden Subgroups
Now, $\Theta$ can be used as a family of problem instances for minimax lower bound of linear bandit. By Theorem 24.1 in [27], we have that with the choice of $\varepsilon=T^{-1/2}$, one has that
$\mathbf{R}_{T}\geq\frac{\exp{-2}}{\sqrt{8}}d\sqrt{T}.$ (16)
@@ Appendix B The Case of Hidden Subgroups with Subexponential Size
B.1 Lower Bound for the Case of Hidden Subgroups with Subexponential Size
See 9
Proof sketch.
First, we show a change of variable between the sparsity and interval partition cases, including the parameter space and set of arms. Next, we demonstrate that mapping the problem instances in the lower bound of sparse bandits in [23, 24] satisfies the constraints in Assumption 6. After that, the proof follows immediately by carrying out step-by-step calculations similar to those in [23, 24].
First, let us define the sparse bandit instances. Let $\mathcal{Z}$ be the set of arm of sparse bandit problem, such that $\forall z\in\mathcal{Z},\;\|z\|_{\infty}\leq 1$. Let $\varphi\in\mathbb{R}^{d}$ be a $d_{0}$-sparse parameters. Now, let us define the mapping that corresponds to change of variable between $\theta$ and $\varphi$ as follows: (1) For $i\in[d-1]$, $\varphi_{i}=\theta_{i}-\theta_{i+1}$, (2) $\varphi_{d}=\theta_{d}$. Therefore, one can write
$\varphi=\underbrace{\begin{bmatrix}1&-1&0&0&\cdots&0&0\\ 0&1&-1&0&\cdots&0&0\\ 0&0&1&-1&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&0&\cdots&1&-1\\ 0&0&0&0&\cdots&0&1\end{bmatrix}}_{W\in\mathbb{R}^{d\times d}}\theta.$ (17)
It is easy to verify that entries of $\theta$ has $k$ equivalent classes if and only if $\varphi$ is $k$ sparse. Now, suppose expected reward for a arm $x$ in interval-partition instance is equal to expected reward for an arm $z$ in sparsity instance, that is,
[…]
Therefore, to guarantee the bijection between sparsity and interval partition instance, we define the set of arm $\mathcal{X}=W^{\top}(\mathcal{Z}):=\left\{W^{\top}z\mid z\in\mathcal{Z}\right\}$.
[…]
$\|\Pi_{m}(x)\|_{2}^{2}$ $=\sum_{k=1}^{\dim(m)}\left(\underbrace{\frac{1}{|S_{k}|}\sum_{i=(\sum_{j<k}|S_{j}|)+1}^{\sum_{j\leq k}|S_{k}|}x_{i}}_{\text{average of entries within class $S_{k}$}}\right)^{2}|S_{k}|$ (19)
$=\sum_{k=1}^{\dim(m)}\frac{1}{|S_{k}|}\left(\sum_{i=(\sum_{j<k}|S_{j}|+1)}^{\sum_{j\leq k}|S_{k}|}x_{i}\right)^{2}.$
Since $x=W^{\top}z$, we have that $x_{i}=-z_{i-1}+z_{i},\;\forall i\in[d]$, with $z_{0}=0$ for convenience. Note that $z_{i}\in[-1,1]\forall i\in[d]$. We have that
$\|\Pi_{m}(W^{\top}z)\|_{2}^{2}$ $=\sum_{k=1}^{\dim(m)}\frac{1}{|S_{k}|}\left(\underbrace{\sum_{i=(\sum_{j<k}|S_{j}|+1)}^{\sum_{j\leq k}|S_{j}|}-z_{i-1}+z_{i}}_{\text{Telescoping sum}}\right)^{2}$ (20)
$\leq\sum_{k=1}^{\dim(m)}\frac{4}{|S_{k}|}\leq 8d_{0}.$
Therefore, we have that, for all $x\in\mathcal{X}=W^{\top}(\mathcal{Z})$, $\left\|\Pi_{m}(x)\right\|_{2}\leq 2\sqrt{2d_{0}}$.
Moreover, we also have that $C_{\min}(\mathcal{X})\geq C_{\min}(\mathcal{Z})$. Particularly, let us denote the exploratory distribution on $\mathcal{Z}$ as
$\rho=\operatorname*{arg\,max}_{\omega\in\Delta(\mathcal{Z})}\lambda_{\min}\left(\mathbb{E}_{z\sim\omega}[zz^{\top}]\right),\quad P:=\mathbb{E}_{z\sim\rho}[zz^{\top}],\quad C_{\min}(\mathcal{Z}):=\lambda_{\min}(P).$ (21)
Define $\mu(x):=\rho((W^{\top})^{-1}z)$, and let $U:=\mathbb{E}_{x\sim\mu}[x^{\top}x]$, we have that $U=W^{\top}PW$. Since all eigenvalue of $W$ is $1$, according to [30], we have that $\lambda_{i}(P)=\lambda_{i}(U)$ for all $i\in[d]$. Therefore, we have that $C_{\min}(\mathcal{X})\geq C_{\min}(\mathcal{Z})$.
Then, using change of variable argument, the calculation of the lower bound is identical to that of [23, 24]. Therefore, we can conclude that all policies must suffer regret
$\mathbf{R}_{T}=\Omega(C_{\min}(\mathcal{X})^{-\frac{1}{3}}d_{0}^{\frac{2}{3}}T^{\frac{2}{3}},\sqrt{dT}).$
B.2 High Probability Prediction Error of Model Selection
@@
Our proof strategy is inspired by [8].
First, we compute the subgaussian norm for the normalised distribution along fews directions $\nu$. For any $x\in\mathcal{X}$, we have that
$\max_{\begin{subarray}{c}\left\|v\right\|_{2}=1,\\ v\in V^{1/2}(\,\overline{\!{m}}),\;\,\overline{\!{m}}\in\,\overline{\!{\mathcal{M}}}\end{subarray}}\left<V^{-1/2}x,v\right>^{2}$ $=\max_{\begin{subarray}{c}\left\|V^{1/2}\omega\right\|_{2}=1,\\ \omega\in\,\overline{\!{m}},\;\,\overline{\!{m}}\in\,\overline{\!{\mathcal{M}}}\end{subarray}}\left<V^{-1/2}x,V^{1/2}\omega\right>^{2}$ (35)
$=\max_{\begin{subarray}{c}\left\|V^{1/2}\omega\right\|_{2}=1,\\ \omega\in\,\overline{\!{m}},\;\,\overline{\!{m}}\in\,\overline{\!{\mathcal{M}}}\end{subarray}}\left<x,\omega\right>^{2}$
$\leq\max_{\,\overline{\!{m}}\in\,\overline{\!{\mathcal{M}}}}\left\|\Pi_{\,\overline{\!{m}}}(x)\right\|_{2}^{2}C_{\min}^{-1}(\mathcal{X})\leq\frac{K_{x}}{C_{\min}(\mathcal{X})}.$
This means that, the distribution $V^{-1/2}x,v\right>^{2}$ has bounded support $[0,\;\frac{K_{x}}{C_{\min}}]$ for any above direction of $v$.
Next, for any $\theta\in\,\overline{\!{m}}$, denote $\theta_{V}:=V^{1/2}\theta$, note that the direction $v$ corresponding to $\theta_{V}$ satisfies (35). Using Hoeffding’s inequality for bounded random variable, there is an absolute constant $c_{1}>0$ such that
$\mathrm{Pr}\left\{\bigg{|}\frac{1}{n}\sum_{i=1}^{n}\left<V^{-1/2}x_{i},\theta_{V}\right>^{2}-\left\|\theta_{V}\right\|_{2}^{2}\bigg{|}\geq\epsilon\left\|\theta_{V}\right\|_{2}^{2}\right\}\leq 2\exp\left(-\frac{nc_{1}C_{\min}^{2}(\mathcal{X})\epsilon^{2}}{K_{x}^{2}}\right)$ (36)
$\iff$ $\mathrm{Pr}\left\{\bigg{|}\frac{1}{n}\left\|XV^{-1/2}\theta_{V}\right\|_{2}^{2}-\left\|\theta_{V}\right\|_{2}^{2}\bigg{|}\geq\epsilon\left\|\theta_{V}\right\|_{2}^{2}\right\}\leq 2\exp\left(-\frac{nc_{1}C_{\min}^{2}(\mathcal{X})\epsilon^{2}}{K_{x}^{2}}\right).$
Let $Z:=\frac{1}{\sqrt{n}}XV^{-1/2}$. From Lemma 5.1 in [5], we know that if the above inequality holds, then
$\left(1-\delta_{\,\overline{\!{M}}}\left(Z\right)\left\|\theta_{V}\right\|_{2}\right)\leq\left\|Z\theta_{V}\right\|_{2}\leq\left(1+\delta_{\,\overline{\!{M}}}\left(Z\right)\left\|\theta_{V}\right\|_{2}\right),$ (37)
holds with probability more than
$1-2\left(\frac{12}{\delta_{\,\overline{\!{M}}}\left(Z\right)}\right)^{2d_{0}}\exp\left(-\frac{c_{1}C_{\min}^{2}(\mathcal{X})n\delta_{\,\overline{\!{M}}}^{2}\left(Z\right)}{K_{x}^{2}}\right),$
for any $\theta_{V}\in V^{1/2}(\,\overline{\!{m}})$ in a subspace $\,\overline{\!{m}}\in\,\overline{\!{M}}$. Take union bound for $\,\overline{\!{M}}$ subspaces, and let
$\delta$ $=2\,\overline{\!{M}}\left(\frac{2}{\delta_{\,\overline{\!{M}}}\left(Z\right)}\right)^{2d_{0}}\exp\left(-\frac{c_{1}C_{\min}^{2}(\mathcal{X})n\delta_{\,\overline{\!{M}}}^{2}\left(Z\right)}{K_{x}^{2}}\right),$
$\iff n$ $=O\left(\frac{K_{x}^{2}}{\delta_{\,\overline{\!{M}}}^{2}\left(Z\right)C_{\min}^{2}(\mathcal{X})}\left(\log(2\,\overline{\!{M}})+d_{0}\log\left(\frac{1}{\delta_{\,\overline{\!{M}}}\left(Z\right)}\right)+\log(\delta^{-1})\right)\right).$
Let $\delta_{\,\overline{\!{M}}}^{2}\left(Z\right)=1/2$. Then, Given $n=O\left(\frac{K_{x}^{2}}{C_{\min}^{2}(\mathcal{X})}\left(\log(2\,\overline{\!{M}})+d_{0}+\log(\delta^{-1})\right)\right)$, for probability at least $1-\delta$, we have that
$\frac{1}{2}\left\|V^{1/2}\theta\right\|_{2}^{2}$ $\leq\frac{1}{n}\left\|XV^{-1/2}V^{1/2}\theta\right\|_{2}^{2},$
$\implies$ $\frac{1}{2}C_{\min}(\mathcal{X})\left\|\theta\right\|_{2}^{2}$ $\leq\frac{1}{n}\left\|X\theta\right\|_{2}^{2}.$
Let $\theta:=\theta_{1}-\theta_{2}$, for any $\theta_{1}\in m$, $\theta_{2}\in m^{\prime}$, for any $m,m^{\prime}\in\mathcal{M}$. We conclude the proof. ∎
B.3 Regret Upper bound
See 12
Proof.
First, let consider Proposition 16, and recall that $\log(M)$ and $\log(\,\overline{\!{M}})$ are both $O(d_{0}\log(d))$. Therefore, for the choice of $t_{1}=\Omega(K_{x}^{2}d_{0}C_{\min}^{-2}(\mathcal{X})\log(d/\delta))$, with probability at least $21-\delta/$ 2, one has that
$C_{\min}(\mathcal{X})\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}^{2}\leq\frac{2}{t_{1}}\left\|X(\theta_{\star}-\widehat{\theta}_{t_{1}})\right\|_{2}^{2}$ (38)
@@ B.3 Regret Upper bound
$\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}\leq c_{2}\sqrt{\frac{\sigma^{2}d_{0}\log\left(\frac{d}{\delta}\right)}{C_{\min}(\mathcal{X})t_{1}}},$ (40)
@@
Let define the pesudo regret as $\widehat{\mathbf{R}}_{T}=\sum_{t=1}^{T}\left<x_{\star}-x_{t},\theta_{\star}\right>$. Denote $\,\overline{\!{m}}\in\,\overline{\!{M}}$ be the subspace contains $\theta_{\star}-\widehat{\theta}_{t_{1}}$. We start by simple regret decomposition as follows.
@@
$\leq R_{\mathrm{\max}}t_{1}+\sum_{t=t_{1}+1}^{T}\left<\theta_{\star}-\widehat{\theta}_{t_{1}},x_{\star}-x_{t}\right>$
$=R_{\mathrm{\max}}t_{1}+\sum_{t=t_{1}+1}^{T}\left<\theta_{\star}-\widehat{\theta}_{t_{1}},\Pi_{\,\overline{\!{m}}}(x_{\star}-x_{t})\right>\;$ $\Big{[}\text{As $\theta_{\star}-\widehat{\theta}_{t_{1}}\in\,\overline{\!{m}}$},\Big{]}$
$\leq R_{\mathrm{\max}}t_{1}+\sum_{t=t_{1}+1}^{T}\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}\left\|\Pi_{\,\overline{\!{m}}}(x_{\star}-x_{t})\right\|_{2}$
$\leq R_{\mathrm{\max}}t_{1}+\sum_{t=t_{1}+1}^{T}2\sqrt{K_{x}}\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2};\;$ $\Big{[}\text{Since $\left\|\Pi_{\,\overline{\!{m}}}(x_{\star}-x_{t})\right\|_{2}\leq 2\sqrt{K_{x}}$ }\Big{]}.$
Now, we invoke Lemma 12. Let $\mathcal{E}$ is the event in that the exploration error is bounded as in Lemma 12, then there is an absolute constant $c_{1}>0$ such that,
$\mathbf{R}_{T}$ $\leq R_{\mathrm{\max}}t_{1}+\mathbb{E}\left[\sum_{t=t_{1}+1}^{T}2\sqrt{K_{x}}\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}\bigg{|}\mathcal{E}\right]+T\mathrm{Pr}(\mathcal{E})R_{\mathrm{\max}}$ (42)
$\leq R_{\mathrm{\max}}t_{1}+c_{1}\sqrt{K_{x}}\sqrt{\frac{\sigma^{2}d_{0}\log(d/\delta)}{C_{\min}(\mathcal{X})t_{1}}}T+T\delta R_{\mathrm{\max}}.$
Let $\delta=1/T$, and $t_{1}=R_{\mathrm{\max}}^{-\frac{2}{3}}\sigma^{\frac{2}{3}}C_{\min}^{-\frac{1}{3}}(\mathcal{X})K_{x}^{\frac{1}{3}}d_{0}^{\frac{1}{3}}T^{\frac{2}{3}}(\log(dT))^{\frac{1}{3}}$, one has that
$\mathbf{R}_{T}=O\left(R_{\mathrm{\max}}^{\frac{1}{3}}\sigma^{\frac{2}{3}}C_{\min}^{-\frac{1}{3}}(\mathcal{X})K_{x}^{\frac{1}{3}}d_{0}^{\frac{1}{3}}T^{\frac{2}{3}}(\log(dT))^{\frac{1}{3}}\right)$ (43)
@@ C.1 Improve Regret Bound with Well-Separated Partitions
Now, by choosing $t_{2}=\Omega\left(C_{\min}^{-2}(\mathcal{X})K_{x}^{2}\log(M\delta^{-1})\right)$, by Proposition 16, we have that with probability at least $31-\delta/$ 3,
$\left\|\bm{f}_{\star}-\Pi_{S_{m}}(\bm{f}_{\star})\right\|_{2}^{2}=\left\|X(\theta_{\star}-\Pi_{S_{m}}(\theta_{\star}))\right\|_{2}^{2}\geq\frac{t_{2}}{2C_{\min}(\mathcal{X})}\left\|\theta_{\star}-\Pi_{S_{m}}(\theta_{\star})\right\|_{2}^{2}\geq\frac{t_{2}\varepsilon_{0}^{2}}{4C_{\min}(\mathcal{X})}$ (49)
@@
$\left\|\bm{f}_{\star}-\bm{\widehat{f}}_{m}\right\|_{2}^{2}+2\left<\bm{\eta},\bm{f}_{\star}-\bm{\widehat{f}}_{m}\right>\geq\frac{t_{2}\varepsilon_{0}^{2}}{8C_{\min}(\mathcal{X})}-5\left(\sigma^{2}d_{0}+\sigma^{2}(\log(3M\delta^{-1}))\right).$ (50)
@@
$\frac{t_{2}\varepsilon_{0}^{2}}{8C_{\min}(\mathcal{X})}-5\left(\sigma^{2}d_{0}+\sigma^{2}(\log(M\delta^{-1}))\right)\geq 0;$ (51)
@@
$t_{2}=\Omega\left(\frac{\sigma^{2}K_{x}^{2}d_{0}\log(d\delta^{-1})}{C_{\min}^{2}(\mathcal{X})\varepsilon_{0}^{2}}\right).$ (52)
The regret upper bound is an immediate consequence of the fact that Algorithm 2 return the true subspace $\hat{m}\ni\theta_{\star}$, combined with the regret bound of OFUL in the exploitation phase and $\delta=1/T$. ∎